# mO Xiusi[#1]

### v1

@@ Abstract
Despite significant advancements in large language models (LLMs), the rapid and frequent integration of small-scale experiences, such as interactions with surrounding objects, remains a substantial challenge. Two critical factors in assimilating these experiences are (1) Efficacy: the ability to accurately remember recent events; (2) Retention: the capacity to recall long-past experiences. Current methods either embed experiences within model parameters using continual learning, model editing, or knowledge distillation techniques, which often struggle with rapid updates and complex interactions, or rely on external storage to achieve long-term retention, thereby increasing storage requirements. In this paper, we propose SELF-PARAM (Self-Updatable Large Language Models with Parameter Integration). SELF-PARAM requires no extra parameters while ensuring near-optimal efficacy and long-term retention. Our method employs a training objective that minimizes the Kullback-Leibler (KL) divergence between the predictions of an original model (with access to contextual information) and a target model (without such access). By generating diverse question-answer pairs related to the knowledge and minimizing the KL divergence across this dataset, we update the target model to internalize the knowledge seamlessly within its parameters. Evaluations on question-answering and conversational recommendation tasks demonstrate that SELF-PARAM significantly outperforms existing methods, even when accounting for non-zero storage requirements. This advancement paves the way for more efficient and scalable integration of experiences in large language models by embedding knowledge directly into model parameters.
@@ 1 Introduction
To achieve long-term Retention, a cognitive system must effectively store its experiences. Existing methods for managing past experiences can be categorized based on whether they require additional modules or parameters beyond the model’s inherent parameters: (1) Methods without Additional Modules or Parameters: These approaches, including continual learning, model editing, and knowledge distillation techniques, embed knowledge directly into the model’s parameters. For instance, model editing methods such as MEND (Mitchell et al., 2022), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), and Model-Editing-FT (Gangadhar & Stratos, 2024) update specific components of the language model to incorporate or modify information. These methods are particularly effective for inserting factual statements like “The highest mountain above sea level is Mount Everest.” However, they may be limited in handling more complex experiences that extend beyond simple facts. Continual learning methods (Sun et al., 2020a; Wu et al., 2024b) can manage more intricate information but typically require compiling a corpus for pre-training or fine-tuning, which may not be suitable for frequent and rapid updates. Moreover, continual learning often relies on Next-Word Prediction (NWP) loss, which may not guarantee strong Efficacy in immediate knowledge integration, and may not be applicable when the experiences to be injected are multi-modal (Wang et al., 2024f; Wu et al., 2024a). Existing Knowledge Distillation methods mainly focus on distilling factual statements (Padmanabhan et al., 2024) or prompts (Choi et al., 2022) into model parameters, which may fall short when injecting complicated experiences. (2) Methods with Additional Modules or Parameters: These approaches employ external components such as text repositories, constructed knowledge bases, or memory modules to store past experiences, thereby facilitating long-term Retention. For example, MemoryLLM (Wang et al., 2024c) and Retrieval-Augmented Generation (Lewis et al., 2020) allow models to access external information during inference. The language model can directly attend to external modules, enhancing Efficacy by providing accurate and relevant information. However, relying on external storage requires maintaining large modules to store all past experiences, thereby increasing storage requirements. Smaller external modules usually offer only marginal benefits (Wang et al., 2024c; Bulatov et al., 2022; 2023).
@@
To address these challenges, we propose a novel method that requires zero additional parameters while yielding nearly optimal Efficacy and maintaining robust long-term Retention. We call this method SELF-PARAM (Self-Evolvable Large language models with Parameter Integration). We formally define context integration as effective when the model can accurately answer questions about specific contexts after incorporating them into its parameters. Based on this definition, we introduce a training objective that minimizes the Kullback-Leibler (KL) divergence between the predictions of an original model (which has access to the context) and a target model (which does not). The intuition behind using KL divergence is that the conditional probabilities of the original model capture the relationships between future queries and the context. If our training sentences are sufficiently diverse (ideally encompassing a wide range of possible queries), the target model should be able to generate appropriate responses to any queries related to the context after training. Since it is infeasible to use all possible sentences for training, we employ an instruct model to generate question-answer pairs about the context, along with sentences sampled from the pre-training dataset, to construct a target sentence set for each context. By minimizing the KL divergence between the conditional distributions of the original model (conditioned on the context) and the predicted distributions of the target model (without the context), we effectively update the target model to internalize the context within its parameters.
@@ 2 Related Work
Methods without extra storage. This category primarily encompasses Model Editing, Continual Learning, and Knowledge Distillation techniques. For model editing methods (Yao et al., 2023), we focus on approaches that directly modify the model parameters, thereby embedding new knowledge within the parameters themselves. MEND (Mitchell et al., 2022) introduces rank-one model edits to update the model parameters, utilizing meta-learning to train a lightweight meta-network for the editing process. Similarly, methods such as ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023) propose closed-form solutions for editing the MLP layers. In contrast, Model-Editing-FT (Gangadhar & Stratos, 2024) fine-tunes the entire model on the factual statements that need to be injected. T-Patcher (Huang et al., 2023) and CaliNet (Dong et al., 2022) store new knowledge in additional neurons. There are also related methods aiming at unlearning knowledge in the language models which might be helpful with enforcing the model to forget some experiences (Wang et al., 2024e). However, these approaches may face challenges related to the continuous growth of parameters when the knowledge to be injected comprises lifespan experiences. As for continual learning, it is usually used in the following situations: (1) Real-Time Assimilation of Dynamic Data: Continual learning facilitates the integration of information from diverse sources such as news (Sun et al., 2020b), scholarly articles Cossu et al. (2022), and social media Cossu et al. (2022). (2) Injection of Extensive Knowledge into Language Models: Jang et al. (2022) presents a method for continual knowledge learning that effectively updates LLMs with new data without compromising previously acquired knowledge. (3) Domain Adaptation: Continual training on new data streams enables models to adapt to specific domains in both language and vision. Cossu et al. (2022) demonstrate this by continuously training models on evolving datasets. Additionally, Ke et al. (2023) propose a soft-masking mechanism to update language models with domain-specific corpora, enhancing performance while preserving general knowledge. Various domain-adapted models have been developed, including those tailored for the financial domain (Xie et al., 2023), E-commerce (Ma et al., 2023), and academic content (Wei et al., 2023). (4) Instruction Tuning for Enhanced Reasoning and Interaction: To enable models to tackle novel tasks, task-incremental continual instruction tuning has been proposed (Wang et al., 2024b; 2023). These tasks encompass mathematical problem-solving (Azerbayev et al., 2023), utilization of calculators, search engines, and databases (Qin et al., 2023). With the rapid development of new tools such as advanced software libraries, novel APIs, and domain-specific utilities (Liang et al., 2023; Jin et al., 2023), there is an increasing necessity to continuously update language models to swiftly adapt and master these innovations. In the realm of Knowledge Distillation, Choi et al. (2022) propose a method for distilling prompts directly into model parameters, effectively replicating scenarios where prompts are appended prior to generation. Later, Padmanabhan et al. (2024) introduces a technique for distilling factual knowledge into model parameters such as “ChatGPT is an AI chatbot developed by OpenAI.” Both approaches demonstrate impressive performance in knowledge injection. However, they may not be directly applicable to the injection of experiences which can be much more complicated.
@@
Methods with extra storage. While the preceding category primarily focus on updating the model itself with new knowledge, another line of research involves storing knowledge in external modules and leveraging these modules to generate responses as needed. This approach encompasses several methodologies: (1) Retrieval-Augmented Generation: This framework involves maintaining a repository of past knowledge, which can be either in the form of knowledge graphs or organized texts (Wang et al., 2024d), from which relevant information is extracted using a retrieval mechanism when generating responses (Gao et al., 2023; Karpukhin et al., 2020; Sarthi et al., 2024; Gutiérrez et al., 2024). (2) Long Context Methods: These methods store all relevant knowledge within the model’s context window. When prompted, the model processes both the extensive context and the input prompt to generate an appropriate response (Wang et al., 2024a; Tworkowski et al., 2023; Chen et al., 2023; Han et al., 2023). (3) Memory-Augmented Methods: This approach typically involves maintaining a memory pool where knowledge is stored. These methods generally include two main operations: (i) Write Operation: Updating the memory pool, often through the compression of textual information into the memory space (Zhou et al., 2023; Wang et al., 2024c). (ii) Read Operation: Retrieving relevant information from the memory pool to inform response generation (Wang et al., 2024c; Zhong et al., 2023). Although memory-augmented methods can achieve impressive performance, they may require significant additional storage space to accommodate complex knowledge externally. In contrast, SELF-PARAM circumvents the need for external storage by embedding all information directly into the model weights via backpropagation. This approach eliminates the additional space requirements associated with external modules, allowing the model to retain and utilize knowledge internally without external dependencies.
@@ 3.2 Construction of Target Sentence Set
To absorb $x$ into $\theta$, we need to construct the target sentence set which contains two parts: (1) sentences related to $x$ (to inject the context $x$) and (2) sentences unrelated to $x$ (to maintain the model’s original abilities). As for (1), we employ an instruct model (we use gpt-4o-mini in our experiments) to generate a set of contextually relevant question-answer pairs as the sentence set. The detailed prompt used for constructing is shown in Appendix A.3. Then for (2), we randomly sample sentences from SlimPajama dataset (Soboleva et al., 2023) as the unrelated set of sentences. With the balanced sentences from the context-related and context-unrelated datasets, we hope to inject the specialized knowledge encapsulated in $x$ into the model while maintaining the model’s broader linguistic capabilities.
@@ 4 Experiments
In our experiments, we mainly consider the following tasks: (1) Single Context Injection (§ 4.1). We adopt SELF-PARAM to inject a single context into the model and then ask the model with the related questions. (2) Batch Context Injection (§ 4.2). We adopt SELF-PARAM to inject multiple contexts simultaneously and then ask the model with the questions related to the injected contexts. (3) Sequential Injection (§ 4.3). We sequentially inject multiple contexts into the model, ensuring that the model does not retain access to earlier contexts during the injection of newer ones. We then test the model’s ability to recall and respond to questions related to the previously injected contexts, thereby assessing its long-term retention. (4) Conversational Recommendation(§ 4.4). We inject up to 1,000 conversations between users and the system into the model to allow the model to read the conversations. After injection, we prompt the model to generate movie recommendations and calculate the recall of these recommendations to measure the recommendation quality.
@@ 4.2 Batch Context Injection
Here we introduce the concept of Storage Complexity $\mathcal{S}$ as proposed in Wang et al. (2024d), which indicates the relative storage space required to store all the contexts that need to be injected compared to the size of all the contexts (such as the total number of tokens in all contexts). As we do not require any additional storage space, SELF-PARAM has $\mathcal{S}$ as zero. In contrast, RAG and long-context methods need to store all contexts, indicating $O(n)$ storage complexity. As MemoryLLM has a fixed-sized memory, the storage complexity is $O(1)$.
@@ 4.3 Sequential Injection
To evaluate the model’s ability to retain knowledge after multiple injections – termed Retention – we conduct sequential injection experiments. Specifically, we aim to determine whether the model can remember previously injected contexts after successive updates. We construct a list of 20 unique contexts from the PwC dataset (Ge et al., 2023) and inject them into the model one after another in a sequential manner. After each injection step, we assess the model’s performance by calculating the QA-F1 score on all questions related to the injected contexts. The QA-F1 scores are averaged across all questions to obtain an average F1-score for each context. To ensure the robustness of our evaluation, we create 50 distinct sequences of contexts. We adopt the Mistral-7B model (Jiang et al., 2023) as our backbone model for these experiments. For each injection step within a sequence, we fine-tune the model using Low-Rank Adaptation (LoRA) weights. After each step, the LoRA weights are merged into the model to incorporate the new context. The results are depicted in Figure 2, illustrating the QA-F1 scores after each sequential injection step.
@@
As shown in Figure 2, our method, SELF-PARAM, demonstrates retention capabilities. Initially, after the first injection, the QA-F1 scores are high on the questions related to the context injected just now. As the number of injections increases to 20 steps, the QA-F1 scores gradually decline to approximately 0.3. Despite this reduction, SELF-PARAM maintains significantly higher performance compared to the base model, which exhibits a QA-F1 score of approximately 0.14 without any injected contexts.

### v2

@@ Abstract
Despite significant advancements in large language models (LLMs), the rapid and frequent integration of small-scale experiences, such as interactions with surrounding objects, remains a substantial challenge. Two critical factors in assimilating these experiences are (1) Efficacy: the ability to accurately remember recent events; (2) Retention: the capacity to recall long-past experiences. Current methods either embed experiences within model parameters using continual learning, model editing, or knowledge distillation techniques, which often struggle with rapid updates and complex interactions, or rely on external storage to achieve long-term retention, thereby increasing storage requirements. In this paper, we propose SELF-PARAM (Self-Updatable Large Language Models by Integrating Context into Model Parameters). SELF-PARAM requires no extra parameters while ensuring near-optimal efficacy and long-term retention. Our method employs a training objective that minimizes the Kullback-Leibler (KL) divergence between the predictions of an original model (with access to contextual information) and a target model (without such access). By generating diverse question-answer pairs related to the knowledge and minimizing the KL divergence across this dataset, we update the target model to internalize the knowledge seamlessly within its parameters. Evaluations on question-answering and conversational recommendation tasks demonstrate that SELF-PARAM significantly outperforms existing methods, even when accounting for non-zero storage requirements. This advancement paves the way for more efficient and scalable integration of experiences in large language models by embedding knowledge directly into model parameters. Code is open-sourced at https://github.com/XinshuangL/SELF-PARAM
@@ 1 Introduction
To achieve long-term Retention, a cognitive system must effectively store its experiences. Existing methods for managing past experiences can be categorized based on whether they require additional modules or parameters beyond the model’s inherent parameters: (1) Methods without Additional Modules or Parameters: These approaches, including continual learning, model editing, and knowledge distillation techniques, embed knowledge directly into the model’s parameters. For instance, model editing methods such as MEND (Mitchell et al., 2022), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), and Model-Editing-FT (Gangadhar & Stratos, 2024) update specific components of the language model to incorporate or modify information. These methods are particularly effective for inserting factual statements like “The highest mountain above sea level is Mount Everest.” However, they may be limited in handling more complex experiences that extend beyond simple facts. Continual learning methods (Sun et al., 2020a; Wu et al., 2024) can manage more intricate information but typically require compiling a corpus for pre-training or fine-tuning, which may not be suitable for frequent and rapid updates. Moreover, continual learning often relies on Next-Word Prediction (NWP) loss, which may not guarantee strong Efficacy in immediate knowledge integration. Existing Knowledge Distillation methods mainly focus on distilling factual statements (Padmanabhan et al., 2024) or prompts (Choi et al., 2022) into model parameters, which may fall short when injecting complicated experiences. (2) Methods with Additional Modules or Parameters: These approaches employ external components such as text repositories, constructed knowledge bases, or memory modules to store past experiences, thereby facilitating long-term Retention. For example, MemoryLLM (Wang et al., 2024c) and Retrieval-Augmented Generation (Lewis et al., 2020) allow models to access external information during inference. The language model can directly attend to external modules, which can enhance Efficacy by providing accurate and relevant information. However, relying on external storage necessitates maintaining large modules to store all past experiences, thereby increasing storage requirements. Smaller external modules usually offer only marginal benefits (Wang et al., 2024c; Bulatov et al., 2022; 2023).
@@
To address these challenges, we propose a novel method that requires zero additional parameters while yielding nearly optimal Efficacy and maintaining robust long-term Retention. We call this method SELF-PARAM (Self-Evolvable Large language models with Parameter Integration). We formally define context injection as effective when the model can accurately answer questions about specific contexts after incorporating them into its parameters. Based on this definition, we introduce a training objective that minimizes the Kullback-Leibler (KL) divergence between the predictions of an original model (which has access to the context) and a target model (which does not). The intuition behind using KL divergence is that the conditional probabilities of the original model capture the relationships between future queries and the context. If our training sentences are sufficiently diverse (ideally encompassing a wide range of possible queries), the target model should be able to generate appropriate responses to any queries related to the context after training. Since it is infeasible to use all possible sentences for training, we employ an instruct model to generate question-answer pairs about the context, along with sentences sampled from the pre-training dataset, to construct a target sentence set for each context. By minimizing the KL divergence between the conditional distributions of the original model (conditioned on the context) and the predicted distributions of the target model (without the context), we effectively update the target model to internalize the context within its parameters.
@@ 2 Related Work
Methods without extra storage. This category primarily encompasses Model Editing, Continual Learning, and Knowledge Distillation techniques. For model editing methods (Yao et al., 2023), we focus on approaches that directly modify the model parameters, thereby embedding new knowledge within the parameters themselves. MEND (Mitchell et al., 2022) introduces rank-one model edits to update the model parameters, utilizing meta-learning to train a lightweight meta-network for the editing process. Similarly, methods such as ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023) propose closed-form solutions for editing the MLP layers. In contrast, Model-Editing-FT (Gangadhar & Stratos, 2024) fine-tunes the entire model on the factual statements that need to be injected. T-Patcher (Huang et al., 2023) and CaliNet (Dong et al., 2022) store new knowledge in additional neurons. However, these approaches may face challenges related to the continuous growth of parameters when the knowledge to be injected comprises lifespan experiences. As for continual learning, it is usually used in the following situations: (1) Real-Time Assimilation of Dynamic Data: Continual learning facilitates the integration of information from diverse sources such as news (Sun et al., 2020b), scholarly articles Cossu et al. (2022), and social media Cossu et al. (2022). (2) Injection of Extensive Knowledge into Language Models: Jang et al. (2022) presents a method for continual knowledge learning that effectively updates LLMs with new data without compromising previously acquired knowledge. (3) Domain Adaptation: Continual training on new data streams enables models to adapt to specific domains in both language and vision. Cossu et al. (2022) demonstrate this by continuously training models on evolving datasets. Additionally, Ke et al. (2023) propose a soft-masking mechanism to update language models with domain-specific corpora, enhancing performance while preserving general knowledge. Various domain-adapted models have been developed, including those tailored for the financial domain (Xie et al., 2023), E-commerce (Ma et al., 2023), and academic content (Wei et al., 2023). (4) Instruction Tuning for Enhanced Reasoning and Interaction: To enable models to tackle novel tasks, task-incremental continual instruction tuning has been proposed (Wang et al., 2024b; 2023). These tasks encompass mathematical problem-solving (Azerbayev et al., 2023), utilization of calculators, search engines, and databases (Qin et al., 2023). With the rapid development of new tools such as advanced software libraries, novel APIs, and domain-specific utilities (Liang et al., 2023; Jin et al., 2023), there is an increasing necessity to continuously update language models to swiftly adapt and master these innovations. In the realm of Knowledge Distillation, Choi et al. (2022) propose a method for distilling prompts directly into model parameters, effectively replicating scenarios where prompts are appended prior to generation. Later, Padmanabhan et al. (2024) introduces a technique for distilling factual knowledge into model parameters such as “ChatGPT is an AI chatbot developed by OpenAI.” Both approaches demonstrate impressive performance in knowledge injection. However, they may not be directly applicable to the injection of experiences which can be much more complicated.
@@
Methods with extra storage. While the preceding category primarily focus on updating the model itself with new knowledge, another line of research involves storing knowledge in external modules and leveraging these modules to generate responses as needed. This approach encompasses several methodologies: (1) Retrieval-Augmented Generation: This framework involves maintaining a repository of past knowledge, which can be either in the form of knowledge graphs or organized texts (Wang et al., 2024d), from which relevant information is extracted using a retrieval mechanism when generating responses (Gao et al., 2023; Karpukhin et al., 2020; Sarthi et al., 2024; Gutiérrez et al., 2024; Izacard et al., 2022; Wang et al., 2022). (2) Long Context Methods: These methods store all relevant knowledge within the model’s context window. When prompted, the model processes both the extensive context and the input prompt to generate an appropriate response (Wang et al., 2024a; Tworkowski et al., 2023; Chen et al., 2023; Han et al., 2023). (3) Memory-Augmented Methods: This approach typically involves maintaining a memory pool where knowledge is stored. These methods generally include two main operations: (i) Write Operation: Updating the memory pool, often through the compression of textual information into the memory space (Zhou et al., 2023; Wang et al., 2024c). (ii) Read Operation: Retrieving relevant information from the memory pool to inform response generation (Wang et al., 2024c; Zhong et al., 2023). Although memory-augmented methods can achieve impressive performance, they may require significant additional storage space to accommodate complex knowledge externally. In contrast, SELF-PARAM circumvents the need for external storage by embedding all information directly into the model weights via backpropagation. This approach eliminates the additional space requirements associated with external modules, allowing the model to retain and utilize knowledge internally without external dependencies.
@@ 3.1.1 The Process of Context Injection
[…]
Moreover, aggregating the KL divergence across all possible sentences $s$ is infeasible. Therefore, we employ sampling techniques to approximate the objective effectively. The sampling methodology for $s$ is detailed in the subsequent section. We denote the set of sentences used to train $\theta^{\prime}$ according to Eq. 3 as $\mathcal{E}$ and we call this set as Target Sentence Set in the remaining paper.
@@ 3.2 Construction of Target Sentence Set
To absorb $x$ into $\theta$, we need to construct the target sentence set which contains two parts: (1) sentences related to $x$ (to inject the context $x$) and (2) sentences unrelated to $x$ (to maintain the model’s original abilities). As for (1), we employ an instruct model (we tried two settings: using gpt-4o-mini and using the corresponding instruct model, i.e., Mistral-7B-Instruct-v0.3 for Mistral-7B and Llama-3-8B-Instruct for Llama-3-8B) to generate a set of contextually relevant question-answer pairs as the sentence set. The detailed prompt used for constructing is shown in Appendix A.3. Then for (2), we randomly sample sentences from SlimPajama dataset (Soboleva et al., 2023) as the unrelated set of sentences. Note that SlimPajama helps preserve the model’s general capabilities, effectively acting as a form of regularization during training. With the balanced sentences from the context-related and context-unrelated datasets, we hope to inject the specialized knowledge encapsulated in $x$ into the model while maintaining the model’s broader linguistic capabilities.
@@ 3.3 Discussion
Where does the performance gain come from?
@@
Analysis of the Computational Cost
We acknowledge that our algorithm could introduce additional computation costs. Compared to the fine-tuning baseline, our approach roughly doubles the overall computational cost in terms of both time and monetary resources. Specifically, our method involves four key steps: (a) generating QA pairs, (b) computing base logits, (c) computing updated logits, and (d) performing backpropagation and optimization. In contrast, continual fine-tuning baselines, such as FT(C) and FT(S) in Tables 1-3, only require steps (c) and (d). Assuming each step is of comparable time complexity, our total cost is approximately twice that of continual fine-tuning. However, given the effectiveness of our method, we argue that these additional costs are justified.
Hallucination Mitigation
As we are encouraging the model to generate answers without seeing the context, there is the risk of hallucination, where the model produces an answer even when it lacks the necessary knowledge. To mitigate this issue, we carefully filter out prompts that require direct access to prior context in the original PwC dataset. For example, prompts such as “summarize the previous context” and “write a title for the above text” were removed, leaving only factual questions. This refined dataset contains context and questions that are based on constant knowledge, such as Question: What language was spoken in the tape recording of Jamal Khashoggi’s murder? Answer: Arabic. These factual statements are less likely to induce hallucination, which helps support the method’s robustness. Regarding the Target Sentence Set, we are not directly fine-tuning the model to predict each word within this set. Instead, we employ KL divergence to transfer the behavior from the original model to the updated one, which may also reduce the risk of hallucination.
@@ 4 Experiments
In our experiments, we mainly consider the following tasks: (1) Single Context Injection (§ 4.1). We adopt SELF-PARAM to inject a single context into the model and then ask the model with the related questions. (2) Batch Context Injection (§ 4.2). We adopt SELF-PARAM to inject multiple contexts simultaneously and then ask the model with the questions related to the injected contexts. (3) Sequential Injection (§ 4.3). We sequentially inject multiple contexts into the model, ensuring that the model does not retain access to earlier contexts during the injection of newer ones. We then test the model’s ability to recall and respond to questions related to the previously injected contexts, thereby assessing its long-term retention. (4) Conversational Recommendation(§ 4.4). We inject up to 1,000 conversations between users and the system into the model to allow the model to read the conversations. After injection, we prompt the model to generate movie recommendations and calculate the recall of these recommendations to measure the recommendation quality. In these tasks, (1), (2), (4) correspond to Efficacy evaluation, while (3) represents the Retention evaluation.
@@ 4.2 Batch Context Injection
Here we introduce the concept of Storage Complexity $\mathcal{S}$ as proposed in Wang et al. (2024d), which indicates the relative storage space required to store all the contexts that need to be injected compared to the size of all the contexts (such as the total number of tokens in all contexts). As we do not require any additional storage space, SELF-PARAM has $\mathcal{S}$ as zero. In contrast, RAG and long-context methods need to store all contexts, indicating $O(n)$ storage complexity. As MemoryLLM has a fixed-sized memory, the storage complexity is $O(1)$. Here for each baseline, we mark the storage complexity to indicate the general complexity which could also be applied to similar methods, rather than the specific complexity for every single baseline, thus offering a more generalized perspective.
@@ 4.3 Sequential Injection
To evaluate the model’s ability to retain knowledge after multiple injections – termed Retention – we conduct sequential injection experiments. Specifically, we aim to determine whether the model can remember previously injected contexts after successive updates. We construct a list of 20 unique contexts from the PwC dataset (Ge et al., 2023) and inject them into the model one after another in a sequential manner. After each injection step, we assess the model’s performance by calculating the QA-F1 score on all questions related to the injected contexts. The QA-F1 scores are averaged across all questions to obtain an average F1-score for each context. To ensure the robustness of our evaluation, we create 50 distinct sequences of contexts. We adopt the Mistral-7B model (Jiang et al., 2023) as our backbone model for these experiments. For each injection step within a sequence, we fine-tune the model using Low-Rank Adaptation (LoRA) weights. After each step, the LoRA weights are merged into the model to incorporate the new context. The results are depicted in Figure 2, illustrating the QA-F1 scores after each sequential injection step. We also include the ablation studies in Appendix B.2
@@
As shown in Figure 2, our method, SELF-PARAM, demonstrates retention capabilities. Initially, after the first injection, the QA-F1 scores are high on the questions related to the context injected just now. As the number of injections increases to 20 steps, the QA-F1 scores gradually decline to approximately 0.3. Despite this reduction, SELF-PARAM maintains significantly higher performance compared to the base model, which exhibits a QA-F1 score of approximately 0.14 without any injected contexts. The diagonal of the figure shows that our model retains its functionality even after multiple rounds of injection, indicating our model’s robustness.